OpenStudy (anonymous):

Audrey has a bag of similar-sized red and green grapes. There are 12 red grapes and 13 green grapes in the bag. If the first 7 grapes she ate were green what is the probability that the 8th grapes she eats is green and the 9th grape she eats is red?

OpenStudy (anonymous):

P(g) * P(r) Probability of it being green at 8th is (13-7)/(12+13 - 7) = 6/18 = 1/3 Probability of it being red at 9th is 12/(12+13 - 8) = 12/17 P(g) * P(r) = 1/3 * 12/17 = 12/51 = 4/17